21 less than 4 times a number is equal to the number. What is the number?
Answer by Flannery(124)

4n-21=n
4n=21+n (add 21)
3n=21 (subtract n)
n=7 (divide by 3)
check
4(7)=21+7
28=28
Q.E.D.